# Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution

### v2

@@ 3.2 Thermal Activations for Super-Resolution
The multi-scale signal representation defined above is a natural match for super-resolution, and particularly well-suited if arbitrary scale factors shall be covered. Still, two challenges must be addressed. First, our formulation restricts the choice of architecture to MLPs with a single hidden layer. Second, while it theoretically guarantees the downsampling operation ($t>t_{0}$), the upsampling operation ($0<t<t_{0}$) remains ill-posed. To narrow down the (infinite) solution space to a unique result, a prior must either be defined in an unsupervised fashion or learned from data.
Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, where $N$ is the number of parameters in a field of the desired size. The hypernetwork maps the low-resolution input image to parameters of local fields, each spanning the area of one pixel of the low-resolution input as originally proposed by LIIF [11] and adopted by most recent ASSR methods [22, 8, 43, 10]. Even though the fields themselves model only a local part of the image, the hypernetwork informs them with rich contextual features collected over a large receptive field. The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [11, 22, 43, 8, 10], followed by one or more layers that produce feature maps of depth $N$. In this work, we use a series of ConvNeXt [26] blocks. The local fields themselves are then supervised with values from high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and time index $t$. The latter ensures that the signal has been filtered to the correct target resolution prior to querying discrete pixel values. Similar to LIIF, the local coordinate system for every field is bounded by $[-0.5,0.5]\times[-0.5,0.5]$. In practice, we directly optimize a single global frequency bank $\textbf{W}_{1}$, rather than having the hypernetwork predict a separate $\textbf{W}_{1}$ for each low-resolution pixel. Not only does this better fit the idea of to represent the signal with a single, consistent basis; it also reduces the total parameter count and empirically leads to better results.
Figure 4: Overview of Thera. A hypernetwork estimates the parameters of local fields. Each field uses thermal activations for multi-scale signal representation and anti-aliasing.
@@
The described scheme, which we call Thera, is depicted in Fig. 4. It allows for arbitrary-scale super-resolution, combining the proposed multi-scale signal representation with the expressivity of proven feature extraction backbones for SR and image restoration. As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors for a whole range of resolutions covered by the training data. E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $\frac{1}{16}<t$. By training on multiple resolutions, we can also make $\kappa$ a trainable parameter that allows the network to adapt to different downsampling operators. Finally, we set the bias terms for the three color channels of every local field $\Phi$ (i.e., $\textbf{b}_{2}$) to the RGB values of the associated low-resolution pixel. Thus, the hypernetwork only predicts field-wise phase shifts $\mathbf{b}_{1}$ and amplitudes $\mathbf{W}_{2}$.
@@
Prior at $\bm{t=0}$. To perform ASSR for out-of-domain scaling factors, we can use an unsupervised prior at $t=0$. Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods. In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well. We use an $\ell^{1}$ variant of TV:
@@
Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically using automatic differentiation tools, rather than falling back to a neighborhood approximation as in most previous work [38].
Implementation and training. We have implemented Thera in JAX [7] to maximize the efficiency of the hypernetwork operations. A PyTorch [36] implementation to fit a Thera field to a single image is also provided. Similar to prior work [11, 22], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$. As corresponding targets, $48^{2}$ random pixels are sampled from the target patch. We train with the Adam optimizer [21] with batch size $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The learning rate is decayed to zero according to a cosine annealing schedule [27]. As reconstruction loss, we minimize the mean absolute error (MAE), to which the TV loss from Eq. 7 is added with a weight of $10^{-4}$. Training data is augmented by standard data augmentations, namely random flipping, rotation, and resizing. As in previous work [42, 24, 43], we implement geometric self-ensembling, where the results for four rotated versions of the input are averaged at test time. We did not observe a performance gain when also including reflections.
@@ 4 Results
We evaluate Thera on a variety of datasets and with three popular feature extractors as backbones. We also conduct an ablation study that supports the design choices described in Sec. 3. Throughout this section, we will evaluate three variants of our method which differ solely in the size of the hypernetwork and the number of field parameters:
@@
  • Thera-S: A tiny version with the number of globally shared components in $\mathbf{W}_{1}$ set to 32, and with only a single layer in hypernetwork, which maps features to field parameters. This version adds only 8,256 parameters on top of the backbone.
  • Thera-M: This version uses 512 shared components and a moderately sized hypernetwork with 5 ConvNeXt layers. Its parameter count of 0.48 M approximately matches that of medium-sized competitors (MetaSR [17], LIIF [11], LTE [22] and CUF [43]).
  • Thera-L: A large version with 512 components and a deeper hypernetwork of 16 ConvNeXt layers, for a total of 1.41 M additional parameters, comparable to CiaoSR [8].
@@
Datasets and metrics. Following previous work, our models are trained with the DIV2K [1] training set, consisting of 800 high-resolution RGB images of diverse scenes. We report evaluation metrics on the official DIV2K validation split (100 images) as well as on a set of standard benchmark datasets: Set5 [6], Set14 [51], BSDS100 [28] (all three containing natural images of low to medium resolution), Urban100 [18] (high-resolution urban scenes), and Manga109 [29] (high-resolution Japanese manga illustrations). Like prior work, we use peak-signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric, where PSNR is computed in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets. Additional metrics are given in Appendix 0.C. Not all numbers could be computed for competing methods, because some do not provide public code or checkpoints, see Appendix 0.E.
@@
Backbones. Following related work, we combine our method with three standard backbones for super-resolution and image restoration: (i) EDSR-baseline [24] (1.22 M parameters); (ii) RDN [52] (22.0 M parameters); and (iii) SwinIR [23] (12.0 M parameters).
Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the DIV2K validation set. The highest PSNR value per backbone and scaling factor is in bold and the second highest is underlined.
Bicubic 31.01 28.22 26.66 24.82 22.27 21.00 20.19 19.59
EDSR- baseline [24] (1.22 M) + interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
MetaSR [17] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
LIIF [11] 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
LTE [22] 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
CUF [43] 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
CiaoSR [8] 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
CLIT [10] 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
Thera-S .008 M 34.75 31.09 29.10 26.84 23.80 22.26 21.26 20.56
Thera-M 0.48 M 34.86 31.21 29.22 26.94 23.87 22.32 21.31 20.60
Thera-L 1.41 M 34.88 31.22 29.24 26.96 23.89 22.33 21.32 20.61
RDN [52] (22.0 M) + interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
MetaSR [17] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
LIIF [11] 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
LTE [22] 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
CUF [43] 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
CiaoSR [8] 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
CLIT [10] 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
Thera-S .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
Thera-M 0.48 M 34.97 31.39 29.43 27.15 24.05 22.49 21.44 20.71
Thera-L 1.41 M 34.98 31.40 29.45 27.16 24.06 22.49 21.45 20.72
SwinIR [23] (12.0 M) + interpol. 35.09 31.42 29.48 27.17 24.08 22.53 21.52 20.82
MetaSR [17] 0.45 M 35.15 31.40 29.33 26.94 23.80 22.26 21.26 20.54
LIIF [11] 0.35 M 35.17 31.46 29.46 27.15 24.02 22.43 21.40 20.67
LTE [22] 0.49 M 35.24 31.50 29.51 27.20 24.09 22.50 21.47 20.73
CUF [43] 0.30 M 35.31 31.56 29.56 27.23 24.10 22.52
CiaoSR [8] 1.43 M 35.28 31.53 29.56 27.26 24.12 22.52 21.48 20.59
CLIT [10] 15.7 M 35.29 31.55 29.55 27.26 24.11 22.51 21.45 20.70
Thera-S .008 M 35.27 31.57 29.58 27.25 24.13 22.55 21.52 20.77
Thera-M 0.48 M 35.29 31.58 29.59 27.27 24.15 22.56 21.52 20.77
Thera-L 1.41 M 35.29 31.59 29.61 27.30 24.17 22.59 21.54 20.79
@@ 4.1 Super-Resolution Performance
We start by evaluating the three variants of our method (Thera-S, -M, and -L) on the hold-out DIV2K validation set, following the setup described above. Table 1 shows the PSNR values obtained by all investigated methods, for both in-distribution (×$\times 2$×$\times 4$) and out-of-distribution (×$\times 6$×$\times 30$) scaling factors. Thera achieves state-of-the-art performance. Thera-L outperforms all previous methods on 21 out of 24 cases, often by a considerable margin (e.g., $\times$3 using EDSR-baseline). Thera-M, which adds 3$\times$ fewer parameters to the backbone, still performs on par with or better than competitors that have comparable parameter counts. For scaling factors higher than ×$\times 2$, Thera-M also matches the performance of the more expensive, attention-based methods CiaoSR and CLIT, despite requiring 30×30$\times$30 × fewer parameters than CLIT.
Notably, even our minimal version Thera-S, with the same anti-aliasing guarantees but with only 8,256 parameters on top of the backbone, reaches very competitive performance. It mostly outperforms all but the most parameter-intensive methods with the EDSR backbone. Using the more powerful SwinIR backbone, Thera-S outperforms all previous methods, including attention-based ones, on 6 out of 8 scaling factors, despite adding almost 2000×2000$\times$2000 × fewer learnable parameters than CLIT (respectively 170×170$\times$170 × fewer than CiaoSR). Similarly, with RDN our smallest variant performs best in 7 of 8 scaling factors. We see this as rather strong evidence for the power of the principled multi-scale observation model that is baked into our network; especially for higher scaling factors.
For a comparison with conventional interpolation, we also report numbers using Lanczos (sinc) resampling with the respective ×$\times 4$ backbone. This baseline is almost always outperformed by dedicated ASSR methods, indicating that the latter do learn scale-specific priors.
Table 2: Evaluation on common benchmark datasets for in-distribution scale factors, with a SwinIR backbone. The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
Method Set5 Set14 B100 Urban100 Manga109
MetaSR 38.26 34.77 32.47 34.14 30.66 28.85 32.39 29.31 27.75 33.29 29.12 26.76 39.46 34.62 31.37
LIIF 38.28 34.87 32.73 34.14 30.75 28.98 32.39 29.34 27.84 33.36 29.33 27.15 39.57 34.68 31.71
LTE 38.33 34.89 32.81 34.25 30.80 29.06 32.44 29.39 27.86 33.50 29.41 27.24 39.63 34.79 31.79
CUF 38.38 34.92 32.83 34.33 30.84 29.05 32.47 29.40 27.88 33.65 29.55 27.32
CiaoSR 38.38 34.91 32.84 34.33 30.82 29.08 32.47 29.42 27.90 33.65 29.52 27.42 39.67 34.84 31.91
CLIT 38.41 34.97 32.86 34.27 30.85 29.08 32.46 29.42 27.91 33.56 29.43 27.25
Thera-S 38.32 34.88 32.75 34.30 30.85 29.07 32.46 29.39 27.88 33.61 29.56 27.25 39.62 34.93 31.91
Thera-M 38.36 34.90 32.80 34.30 30.86 29.12 32.47 29.40 27.89 33.58 29.56 27.27 39.70 35.02 32.04
Thera-L 38.41 34.93 32.82 34.33 30.87 29.11 32.49 29.42 27.91 33.67 29.62 27.33 39.71 35.01 32.07
We further report the performance of Thera on five popular benchmark datasets in Tab. 2. For these experiments, we use the SwinIR backbone. Overall, due to the unreasonably small data volume and out-of-distribution nature of some legacy datasets (e.g., 5 small images for Set5), the results are less consistent. We posit that future work on ASSR would benefit from switching to a larger, statistically more meaningful test bed. Nonetheless, our methods perform best in 12 out of 15 cases. We once again observe that the performance of Thera-S is often comparable to that of methods with orders of magnitude higher parameter counts.
When visually inspecting the super-resolution results – for examples see Fig. 5 – we observe that Thera achieves a more faithful and perceptually more convincing reconstruction, particularly in the presence of repeating structures. This again highlights the advantage of hard-wiring a theoretically principled sampling model that rules out signal aliasing. For further qualitative results see Appendix 0.B.
@@
LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
Figure 5: Qualitative examples from four datasets for $\times$6 super-resolution. We use the SwinIR [23] backbone for all methods. Differences between methods are highlighted with bounding boxes. More examples are given in Appendix 0.B.
4.2 Multi-Scale Consistency
We are interested in learning a super-resolution model that is consistent across scales. For instance, if we super-resolve an image with a scaling factor $\times$4 then downsample it by a factor $\times$2, we expect the result to match what we get by directly super-resolving the original image by a factor $\times$2. We report in Tab. 3 the quality of this match for several pairs of scaling factors, both for Thera with its in-built observation model and for three comparison methods where the observation model must be discovered from data. First, we notice that higher super-resolution quality does not correlate with better scale consistency, e.g., LTE in most cases achieves better consistency than CiaoSR. Second, while other methods generally exhibit good consistency, even for out-of-distribution (OOD) scales, our formal model is superior for all scaling factor combinations, often by a large margin, while requiring only a single parameter $\kappa$ to be learned (if any, since $\kappa$ can also be fixed as we show below).
Table 3: Multi-scale consistency (in PSNR) of super-resolution methods across various pairs of scaling factors.
$(2,4)$ $(2,6)$ $(2,12)$ $(2,24)$ $(3,12)$ $(3,24)$ $(4,24)$
LIIF [11] 44.88 43.75 42.21 41.48 43.64 42.52 44.20
LTE [22] 44.10 44.40 44.37 44.40 48.33 48.48 51.42
CiaoSR [8] 44.24 43.31 42.10 41.56 41.47 41.03 45.47
Thera-S 45.79 46.67 46.80 46.89 52.38 52.66 54.84
Thera-M 45.26 46.10 46.34 46.41 51.91 52.16 54.64
Thera-L 45.42 46.24 46.43 46.47 51.56 51.77 54.33

### v3

@@ 3.2 Learning a Super-Resolution Prior
The multi-scale signal representation inherent in neural heat fields is a natural match for ASR. Still, two challenges must be addressed. First, our formulation restricts the choice of architecture to MLPs with a single hidden layer. Second, while it theoretically guarantees the downsampling operation ($t>1$), the upsampling operation ($10<t<$ 1) remains ill-posed. To narrow down the (infinite) solution space to a unique result, a prior must either be defined in an unsupervised fashion or learned from data. Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, mapping the low-resolution input image to the $N$ parameters of each local neural heat field. As originally proposed by LIIF [10] and adopted by recent ASR methods [24, 7, 44, 9], each local field spans the area of one pixel of the low-resolution input. It is important that even though the fields themselves model only a local part of the image, the hypernetwork informs them with contextual features collected over a large receptive field. The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [10, 24, 44, 7, 9], followed by one or more layers that produce feature maps of depth $N$.
During training, the local fields are supervised with values of high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and time index $t$ (ensuring that the signal is correctly blurred for the target resolution), and the entire architecture is optimized end-to-end. In practice, we directly optimize a single global frequency bank $\textbf{W}_{1}$, rather than having the hypernetwork predict a separate $\textbf{W}_{1}$ for each low-resolution pixel. Not only does this better fit the idea to represent the signal with a single, consistent basis, it also reduces the total parameter count.
@@
The described scheme, which we call Thera, is depicted in Fig. 3. It allows for arbitrary-scale super-resolution, combining the multi-scale signal representation within neural heat fields with the expressivity of proven feature extraction backbones for SR and image restoration. As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors for a whole range of resolutions covered by the training data. E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $t>\frac{1}{16}$. By training on multiple resolutions, we can also make $\kappa$ a trainable parameter that allows the network to adapt to different downsampling operators. Finally, we set the bias terms for the three color channels of every local field $\Phi$ (i.e., $\textbf{b}_{2}$) to the RGB values of the associated low-resolution pixel. Thus, the hypernetwork only predicts field-wise phase shifts $\mathbf{b}_{1}$ and amplitudes $\mathbf{W}_{2}$.
3.3 Total Variation at $\bm{t=0}$
@@
To allow Thera to better generalize to higher, out-of-domain scaling factors, we can place an unsupervised regularizer at $t=0$. Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods. In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well [11]. We use an $\ell^{1}$ variant of TV,
@@
Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically by automatic differentiation, rather than falling back to a neighborhood approximation as in most previous work [38]. We further motivate this approach in Fig. 5, which demonstrates that our method faithfully recovers the gradients of super-resolved images.
3.4 Implementation and Training
Thera is implemented in JAX [6]. Similar to prior work [10, 24, 7, 56], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$. As corresponding targets, $48^{2}$ random pixels are sampled from the target patch. We train with standard augmentations (random flipping, rotation, and resizing), using the Adam optimizer [23] with a batch size of $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The learning rate is decayed to zero according to a cosine annealing schedule [30]. We use MAE as reconstruction loss, to which the TV loss from Eq. 5 is added with a weight of $10^{-4}$. Like previous work [43, 26, 44], we employ geometric self-ensembling (GSE) instead of the local self-ensembling introduced in LIIF [10]. In GSE, the results for four rotated versions of the input are averaged at test time. Including reflections did not improve performance.
@@ 4 Results
Throughout this section we evaluate three variants of our method, which differ solely in the size of the hypernetwork and the number of field parameters:
@@
  • Thera Air: A tiny version with the number of globally shared components in $\mathbf{W}_{1}$ set to 32, and the hypernetwork being a single $1\times 1$ convolution that maps features to field parameters. This version adds only 8,256 parameters on top of the backbone.
  • Thera Plus: A balanced version that employs an efficient ConvNeXt-based [29] hypernetwork. Its parameter count of $\approx$1.41 M matches that of recent medium-sized competitors like [7].
  • Thera Pro: The strongest version uses a high-capacity, transformer-based [28, 25] hypernetwork. Its added parameter count is $\approx$4.63 M, still less than the most recent competitor [56] and much smaller than [9].
@@
Datasets and metrics. Following previous work, our models are trained with the DIV2K [1] training set, consisting of 800 high-resolution RGB images of diverse scenes. We report evaluation metrics on the official DIV2K validation split as well as on standard benchmark datasets: Set5 [5], Set14 [53], BSDS100 [31], Urban100 [19], and Manga109 [32]. Following prior work, we use peak signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric and compute it in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets. Additional quantitative results are given in Appendix C. Not all numbers could be computed for competing methods for which code or checkpoints were not publicly shared (see Appendix F).
@@
Backbones. Like previous work, we combine our method with two standard backbones for super-resolution and image restoration: (i) EDSR-baseline [26] (1.22 M parameters) and (ii) RDN [54] (22.0 M parameters).
@@ 4.1 Super-Resolution Performance
Low-res input LIIF [10] SRNO [45] MSIT [56] Thera Pro (ours) GT
DIV2K
@@
Manga109
Figure 4: Qualitative examples for a representative $\times$6 scale factor, with an RDN [54] backbone for all methods. Best viewed zoomed in.
Qualitative results. Upon visual inspection – see Fig. 4 for examples – we observe that Thera produces results that are both perceptually convincing and more correct, particularly in the presence of repeating structures. Neural heat fields enable Thera to reproduce a high level of detail without suffering from aliasing, no matter the sampling scale (see also Fig. 9 in Appendix B).
Bicubic 31.01 28.22 26.66 24.82 22.27 21.00 20.19 19.59
EDSR- baseline [26] (1.22 M) + Sinc interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
MetaSR [18] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
LIIF [10] 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
LTE [24] 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
CUF [44] 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
CiaoSR [7] 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
CLIT [9] 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
SRNO [45] 0.80 M 34.85 31.11 29.16 26.90 23.84 22.29 21.27 20.56
MSIT [56] 4.83 M 34.95 31.23 29.22 26.94 23.83 22.27 21.26 20.54
Thera Air (ours) .008 M 34.75 31.09 29.10 26.84 23.80 22.26 21.26 20.56
Thera Plus (ours) 1.41 M 34.89 31.22 29.24 26.96 23.89 22.34 21.32 20.61
Thera Pro (ours) 4.63 M 35.19 31.50 29.51 27.19 24.09 22.51 21.48 20.73
RDN [54] (22.0 M) + Sinc interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
MetaSR [18] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
LIIF [10] 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
LTE [24] 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
CUF [44] 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
CiaoSR [7] 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
CLIT [9] 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
SRNO [45] 0.80 M 35.16 31.42 29.42 27.12 24.03 22.46 21.41 20.68
MSIT [56] 4.83 M 35.16 31.42 29.42 27.11 23.99 22.42 21.37 20.65
Thera Air .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
Thera Plus 1.41 M 35.00 31.40 29.44 27.16 24.06 22.49 21.45 20.71
Thera Pro 4.63 M 35.25 31.56 29.57 27.25 24.14 22.56 21.52 20.77
Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the held-out DIV2K validation set. The highest PSNR value per backbone and scaling factor is bold and the second highest is underlined.
Quantitative results. We first evaluate the three variants of our method on the held-out DIV2K validation set, following the setup described above. Table 1 shows PSNR values for all tested methods, for both in-distribution (×$\times 2$ to ×$\times 4$) and out-of-distribution (×$\times 6$ to ×$\times 30$) scaling factors. Thera Pro outperforms all competing methods at all scaling factors, often by a substantial margin (e.g., 29.51 vs. 29.22 on EDSR ×$\times 4$), even though its parameter overhead on top of the backbone is lower compared to the second-best method MSIT [56], and less than a third of CLIT [9]. Interestingly, even our minimal variant Thera Air – with only about 8000 parameters on top of the backbone – performs on par with or better than methods of much higher parameter count. This supports our claim that hard-wiring a theoretically principled sampling model, which rules out signal aliasing, enables better generalization and higher-fidelity reconstruction. For comparison with conventional interpolation, we also report numbers obtained with Lanczos (sinc) resampling on top of the respective ×$\times 4$ backbone. This baseline is consistently outperformed by dedicated ASR methods, indicating that the latter do learn scale-specific priors.
Method Set5 Set14 B100 Urban100 Manga109
MetaSR 38.22 34.63 32.38 33.98 30.54 28.78 32.33 29.26 27.71 32.92 28.82 26.55
LIIF 38.17 34.68 32.50 33.97 30.53 28.80 32.32 29.26 27.74 32.87 28.82 26.68 39.26 34.21 31.20
LTE 38.23 34.72 32.61 34.09 30.58 28.88 32.36 29.30 27.77 33.04 28.97 26.81 39.28 34.32 31.30
CUF 38.28 34.80 32.63 34.08 30.65 28.92 32.39 29.33 27.80 33.16 29.05 26.87
CiaoSR 38.29 34.85 32.66 34.22 30.65 28.93 32.41 29.34 27.83 33.30 29.17 27.11 39.51 34.57 31.57
CLIT 38.26 34.80 32.69 34.21 30.66 28.98 32.39 29.34 27.82 33.13 29.04 26.91
SRNO 38.32 34.84 32.69 34.27 30.71 28.97 32.43 29.37 27.83 33.33 29.14 26.98 39.52 34.67 31.61
MSIT 38.31 34.85 32.72 34.26 30.70 28.97 32.42 29.35 27.81 33.27 29.14 26.93 39.44 34.62 31.58
Thera Air 38.18 34.75 32.60 34.13 30.70 28.95 32.33 29.29 27.80 33.16 29.14 26.91 39.03 34.57 31.66
Thera Plus 38.11 34.67 32.56 34.20 30.67 28.97 32.26 29.28 27.81 33.14 29.15 26.97 38.69 34.42 31.65
Thera Pro 38.36 34.88 32.79 34.43 30.85 29.08 32.46 29.39 27.87 33.63 29.58 27.26 39.62 34.98 31.98
Table 2: Results on common benchmark datasets for in-distribution scale factors with an RDN [54] backbone. The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
Like earlier work, we further report the performance of Thera on five popular benchmark datasets with an RDN backbone in Tab. 2. Our method again outperforms all competing methods in all settings, often substantially (e.g., 29.58 vs. 29.14 on Urban100 ×$\times 3$). We hypothesize that Thera’s hard-wired PSF is also beneficial when generalizing to unseen datasets. Once again we observe that the performance of Thera Air is often comparable to that of methods with orders of magnitude higher parameter counts.
Figure 5: Thera reconstructs a signal $\Phi$ and its gradient $\nabla_{\mathbf{x}}\Phi$ more faithfully than a ReLU-based competitor [10]. Due to its natural, Fourier-inspired representation, Thera is also infinitely differentiable, while ReLU-based competitors approximate the signal as a piecewise-linear function with null higher derivatives (last row).
Fidelity of the Signal and its Derivatives. Neural fields with periodic activation functions have been shown to be superior when it comes to fitting high-resolution, natural signals, and to correctly recovering their derivatives [41]. We observe similar effects for Thera, whose thermal activations at $t=0$ can be seen as a special case of periodic activations, cf. Figure 5. In fact, due to the use of thermal activations – and unlike all prior work based on multi-layer ReLU-activated fields – Thera is infinitely differentiable.